OpenStudy (anonymous):

PLEASE HELP ME 1. What is the slope of the line that passes through the points (3, 1) and (8, 3)? A. 2/5 B. 5/2 C. -2/5 D. -5/2

OpenStudy (anonymous):

do you have any idea how to find a slope ?

OpenStudy (anonymous):

It would be A.2/5 because with Rise over Run you go 2 up and to the right 5

OpenStudy (anonymous):

can you help me with this Based on the two previous questions, find the line that is perpendicular to the line 6x - 3y = 3 and passes through the point (4, -2). You may write your final answer in any form (Point-Slope, Slope-Intercept, or Standard Form

OpenStudy (anonymous):

y=6/3x+1 I think. @iGreen am i correct?

OpenStudy (igreen):

Try not to give answers @GhostVortex , we're supposed to guide them to the answer, so we need their participation :)

OpenStudy (anonymous):

ok, sorry @iGreen

OpenStudy (anonymous):

thank you so much @GhostVortex

OpenStudy (astrophysics):

Just use the slope formula \[m = \frac{ y_{2}-y _{1} }{ x _{2}-x _{1} }\]

OpenStudy (anonymous):

thats the point , i could've gave the answer , but after @iGreen's advice , i'm just guiding B)

OpenStudy (tkhunny):

@iGreen has it right. Guide, Search. Don't just dump stuff.